# Apparently ultra-long period radio sources from self-lensed pulsar-black hole binaries

### v1

@@ Abstract
Pulsar-black hole (BH) close binary systems, which have not been found yet, are unique laboratories for testing theories of gravity and understanding the formation channels of gravitational-wave sources. We study the self-gravitational lensing effect in a pulsar-BH system on the pulsar’s emission. Because this effect occurs once per orbital period for almost edge-on binaries, we find that it could generate apparently ultra-long period (minutes to hours) radio signals when the intrinsic pulsar signal is too weak to detect. Each of such lensed signals, or ’pulse’, is composed of a number of amplified intrinsic pulsar pulses. The model is applied to three recently found puzzling long-period radio sources: GLEAM-X J1627, PSR J0901-4046, and GPM J1839–10. To explain their observed signal durations and periods, the masses of their lensing components would be $\sim 10^{4}\,\rm M_{\odot}$, $\sim 4\,\rm M_{\odot}$ and $10^{3-6}\,\rm M_{\odot}$, respectively. Their binary coalescence times are from a few tens to thousands of years. We estimate that a radio telescope with a sensitivity of $10\,\rm mJy$ could detect approximately $20$ systems that emit such signals in our galaxy. For a binary containing a millisecond pulsar and a stellar-mass BH, the Shapiro delay effect would cause a $\geq 10\%$ variation of the profile width for the sub-pulses in such lensed signals.
@@ 1 Introduction
Radio pulsars, after their discovery in 1967 (Hewish et al., 1968), were soon identified as rapidly rotating and highly magnetized neutron stars (NS). Their observed periods of signals are short and stable, making them one of the most precise clocks in the universe. As a result, pulsar binaries become unique laboratories to study stellar evolution (Ferdman et al., 2020; Stairs, 2004), NS magnetospheric physics (Breton et al., 2012; Lomiashvili & Lyutikov, 2014) and the strong gravitation near compact objects (Kramer et al., 2006; Ransom et al., 2014). Up to now, more than hundreds of pulsar binary systems have been found. However, among them, few binaries harbour another compact object besides pulsar. For instance, only fifteen double NS systems were detected, and thirteen of them were found to contain millisecond pulsars (Pol et al., 2019). Moreover, the pulsar-black hole (BH) binaries are even considered the holy grail of pulsar astronomy, but their only observed evidence so far is the gravitational-wave (GW) signals from two NS-BH merge events detected by the LIGO–Virgo detector network (Abbott et al., 2021).
@@
This picture is partly motivated by three recently found puzzling radio sources: GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GBM J1839–10 (Hurley-Walker et al., 2023). Among them, PSR J0901-4046 (Caleb et al., 2022) shows variable pulse profiles and has the shortest period $\sim 1.3\,\rm min$, which is still much longer compared to a normal pulsar. As for GLEAM-X J1627 (Hurley-Walker et al., 2022), its $\sim 18.2\,\text{min}$ period and multiple-peak pulse structure are similar to the observational appearance of a self-lensed pulsar-BH binary that our model predicts. GBM J1839–10 (Hurley-Walker et al., 2023) is also characterized by an ultra-long period of $\sim 22\,\text{min}$. Moreover, it has been found to have been actively existing for three decades, which makes this source even more mysterious.
This paper is organized as follows. In §2, we describe the toy model for the self-lensing and the assumed intensity profile of the pulsar’s intrinsic emission. In §3, we predict the appearance of ultra-long period radio signals in an edge-on pulsar-BH binary and apply the model to those recently founded puzzling radio sources. We estimate the number of observable self-lensed pulsar-BH binaries in §4. Additionally, we consider the Shapiro delay effect and estimate its resulting distortion of the lensed pulse profiles in §5. Finally, we summarize the results in §6.
@@ 2.2 Pulsar emission and self-lensing signals
With different parameter values, the lensed signals show diverse appearances. In Fig. 3, panels (a) and (c) exhibit the cases in which $M_{\rm BH}=10\rm M_{\odot}$. In comparison to (a), the sub-pulses in (c) have a higher level of $F_{0}$ and $\phi$, which makes the amplified $F_{0}$ comparable to the peak fluxes of the surrounding lensed sub-pulses and results in a larger duty cycle. For the cases in panels (b), (d) and (e), the BH mass is $20\rm M_{\odot}$. Among them, panel (b) showcases a situation where the pulsar spins faster than in (d) and (e), which enables more sub-pulses to be detected in a self-lensing signal. Therefore, its overall shape of the signal resembles the amplification factor evolution shown in Fig. 2. The orbital inclination is adjusted to higher ($cos\,i=3\times 10^{-4}$) and lower ($cos\,i=6\times 10^{-6}$) levels in panels (d) and (e), corresponding to flatter or steeper signal profiles, respectively. In addition, we present an extreme case in panel (f), where the BH mass $M_{\rm BH}=2\times 10^{4}\rm M_{\odot}$, the spin period $P=4\rm s$ and the orbital period $P_{\rm b}=0.33\rm hr$.
@@ 3.1 Puzzling radio sources
When carrying out this work, we came across three recently discovered interesting radio sources, namely GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839–10 (Hurley-Walker et al., 2023). They all exhibit characteristics similar to extremely slow-rotating pulsars with spin periods of approximately $18.2\,\rm min$, $1.3\,\rm min$, and $22.0\,\rm min$, respectively.
@@
In contrast, GLEAM-X J1627 (Hurley-Walker et al., 2022) showcases signals with a much longer period ($18.2\,\rm min$) and a relatively stable pulse morphology, which exhibits a multi-peak structure. Its pulse duration ranges from $30$ to $60\,\rm s$, and the period of $18.18\,\rm min$ poses challenges to existing pulsar models. Even more puzzling, after three months of activity, it disappeared from observations and could not be found in available data. This ultra-long period but short-lived behaviour has prompted a range of interpretations, including a WD with relatively strong magnetic fields (Katz, 2022; Loeb & Maoz, 2022) and a spinning-down magnetar (Suvorov & Melatos, 2023; Tong, 2023).
@@
Therefore, we apply our model to these sources. In Fig. 4, PSR J0901-4046 is represented just by a blue dot due to its well-constrained pulse duration. For GLEAM-X J1627 and GPM J1839–10, we plot them as blue dots with error bars, where the dots correspond to the median value of their pulse durations and the bars represent a range between the maximum and minimum values. The results suggest BH masses of $4\,\rm M_{\odot}$ and $\sim 10^{4}\,\rm M_{\odot}$ for PSR J0901-4046 and GLEAM-X J1627, with corresponding coalescence times of $10^{2}\,\rm yrs$ and $\sim 6\times 10^{2}\,\rm yrs$, respectively. For GPM J1839–10, its variable signal duration makes the results range in several orders of magnitudes: The BH mass ranges from $\sim 10^{3}$ to $\sim 10^{6}\,\rm M_{\odot}$, with coalescence times spanning from $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$.
@@
In our estimation, both GLEAM-X J1627 and GPM J1839–10 seem to harbour an intermediate-mass BH. However, the coalescence time derived for GLEAM-X J1627 appears excessively long, inconsistent with its reported three-month active period. Conversely, PSR J0901-4046 has been staying active since its detection, which aligns well with the estimated 100-year lifetime. However, the derived mass for its lens is $\sim 4\,\rm M_{\odot}$, which lies in the $2-5\,\rm M_{\odot}$ mass gap and is not typical for observed stellar-mass BHs so far. Regarding GPM J1839–10, the derived coalescence time range ($\sim 30$ to $\sim 3\times 10^{3}\,yrs$) aligns with its three-decade activity, implying monitoring a continuing activity of the system for at least another 30 years, which can be verified by the future monitoring.
@@
Our model predicts negative, large period derivatives ($\sim 10^{-8}\,\rm s\,s^{-1}$), apparently inconsistent with the reported $\dot{P}$ of the three sources, especially for PSR J0901-4046: It has a positive $\dot{P}\approx 2.25\times 10^{-13}\,\rm s\,s^{-1}$ (Caleb et al., 2022). The other two sources, GLEAM-X J1627 and GPM J1839–10, have reported $\dot{P}$ below two upper limits: $\leq 1.2\times 10^{-9}\,\rm s\,s^{-1}$ (Hurley-Walker et al., 2022) and $\leq 3.6\times 10^{-13}\,\rm s\,s^{-1}$ (Hurley-Walker et al., 2023).
4 The number estimation for detection
Here we estimate the number of observable self-lensed pulsar-BH systems within the Milky Way. We first compute the number of galactic NS-BH binaries that can reach a threshold lensing amplification $A_{\rm th}$ in §4.1. This calculation also incorporates constraints from recent NS-BH merger events that are detected by LIGO-Virgo. Then in §4.2, we estimate the number of galactic self-lensed pulsar-BH binaries assuming a uniform radio luminosity for all pulsars.
4.1 The binary number estimation for a threshold amplification
$\tau=\frac{\pi R_{\rm E}^{2}}{4\pi a^{2}}\,=\,\frac{GM_{\rm BH}}{ac^{2}}.$ (12)
It is also the possibility for $u_{\rm min}\leq 1$, which corresponds to that the closest angular separation between the pulsar and BH is smaller than the angular Einstein radius. In observation, a threshold amplification $A_{\rm th}$ is needed to make the lensed pulse flux exceed the telescope’s sensitivity and become detectable (i.e. as long as $A_{\rm max}>A_{\rm th}$, signals composed of self-lensed pulses can be observed). This implies: $u_{\rm min}\leq u_{\rm th}=A_{\rm th}^{-1}$. The possibility that this constraint holds is:
$\tau(A_{\rm th})$ $=\,\frac{\pi(u_{\rm th}R_{\rm E})^{2}}{4\pi a^{2}}$ (13)
$p(M_{\rm BH})$ $\equiv\frac{1}{N_{\rm tot}}\frac{dN}{dM_{\rm BH}}$ (14)
where $t_{\rm coal}(P_{\rm b},M_{\rm BH})$ is the coalescence time given by Eq. (11), and $\mathcal{R}_{i}$ is the merger rate, which varies depending on the binary type. Because the binary studied here is the NS-BH binary, $\mathcal{R}_{\rm i}=45\,\rm Gpc^{-3}\,yr^{-1}$ is adopted, based on the two NS-BH coalescence events detected by LIGO-Virgo (Abbott et al., 2021).
[…]
We set $M_{\rm NS}=1.4\,\rm M_{\odot}$ and allow $M_{\rm BH}$ to range from $M_{\rm BH,min}=10\,\rm M_{\odot}$ to $M_{\rm BH,max}=100\,\rm M_{\odot}$. To identify such systems, a sufficient number of signals must be observed. Therefore, we choose $P_{\rm min}=30s$. This ensures that the binaries considered could all exist at least $\sim 1\rm yr$ until coalescence. Because the fraction of binaries with a short period ($\sim\rm min$) is small, $P_{\rm max}$ is set to be $\gg P_{\rm min}$. Thus Eq. (16) can be written as:
@@ 4.1 The binary number estimation for a threshold amplification
$\mathcal{N}_{\rm lens}(A_{\rm th})\approx 80\times\left(\frac{P_{\rm max}}{1\rm day}\right)^{2}\left(\frac{A_{\rm th}}{10}\right)^{-2}.$ (17)
This indicates that within the Milky Way, approximately eighty NS-BH binaries could satisfy the threshold amplification of $A_{\rm th}=10$.
4.2 The number of observable self-lensed pulsar-BH binaries
$f_{\rm c}=\frac{\Omega_{\rm c}}{4\pi}=\frac{1}{4}\,\sin^{2}(\phi).$ (18)
$\phi=\pi/12$ is set hereafter.
[…]
$\approx 5\times\left(\frac{P_{\rm max}}{1\,\rm day}\right)^{2}\left(\frac{A_{\rm th}}{2}\right)^{-2}\left(\frac{F_{\rm lim}}{10\,\rm mJy}\right)^{-1}\left(\frac{L_{\nu}}{1\,\rm Jy\,kpc^{2}}\right).$
Using $\rho(P_{\rm b},M)$ in a distance interval $dD$:
the number of the ultra-long period radio sources is:
[…]
$\approx 16\times\left(\frac{P_{\rm max}}{1\,\rm day}\right)^{2}\left(\frac{F_{\rm lim}}{10\,\rm mJy}\right)^{-1}\left(\frac{L_{\nu}}{1\,\rm Jy\,kpc^{2}}\right).$
In Figure 5, we explore the correlation between the estimated numbers ($\mathcal{N}_{\rm mod}$ and $\mathcal{N}_{\rm ulp}$) and different values of $P_{\rm max}$, $L_{\nu}$ and $F_{\rm lim}$. Notably, when $F_{\rm lim}$ is at a moderate level ($\gtrsim 10^{-4}\,\rm mJy$ for black lines or $\gtrsim 10^{-3}\,\rm mJy$ for red lines), these two numbers can be effectively described by the power-law function of the free parameters, as shown in the second lines of Eqs. (21) and (24). However, when the sensitivity reaches a level at which the telescope can detect the intrinsic pulsar emission from every pulsar-BH binary in the Milky Way, the observed signals will manifest as modulated pulsar signals, and no puzzling ultra-long period sources can be found. Consequently, $\mathcal{N}_{\rm mod}$ becomes steady and $\mathcal{N}_{\rm ulp}$ sharply drops.
As Eqs. (21) and (24) show, if all the pulsar’s radio luminosity is $1\,\rm Jy\,kpc^{2}$, there will be approximately $20$ pulsar-BH binaries in the Milky Way with observable self-lensing signals. Their periods range from hours to a day and can be detected by a telescope with a sensitivity of $10\,\rm mJy$. Among these systems, $16$ binaries would exhibit ultra-long period characteristics and the others are only modulated pulsar signals.

### v2

@@ Abstract
Pulsar-black hole (BH) close binary systems, which have not been found yet, are unique laboratories for testing theories of gravity and understanding the formation channels of gravitational-wave sources. We study the self-gravitational lensing effect in a pulsar-BH system on the pulsar’s emission. Because this effect occurs once per orbital period for almost edge-on binaries, we find that it could generate apparently ultra-long period (minutes to hours) radio signals when the intrinsic pulsar signal is too weak to detect. Each of such lensed signals, or ’pulse’, is composed of a number of amplified intrinsic pulsar pulses. We estimate that a radio telescope with a sensitivity of $10\,\rm mJy$ could detect $\sim$ a few systems that emit such signals in our galaxy. The model is applied to three recently found puzzling long-period radio sources: GLEAM-X J1627, PSR J0901-4046, and GPM J1839-10. To explain their observed signal durations and periods, the masses of their lensing components would be $\sim 10^{4}\,\rm M_{\odot}$, $\sim 4\,\rm M_{\odot}$ and $10^{3-6}\,\rm M_{\odot}$, respectively, with their binary coalescence times ranging from a few tens to thousands of years. However, the implied merger rates (as high as $\sim 10^{3-4}\,\rm Myr^{-1}$ per galaxy) and the large period decay rates ($>10^{-8}\,\rm s\,s^{-1}$) tend to disfavour this self-lensing scenario for these three sources. Finally, for a binary containing a millisecond pulsar and a stellar-mass BH, the Shapiro delay effect would cause a $\geq 10\%$ variation of the profile width for the sub-pulses in such lensed signals.
@@ 1 Introduction
Radio pulsars, after their discovery in 1967 (Hewish et al., 1968), were soon identified as rapidly rotating and highly magnetized neutron stars (NS). Their observed periods of signals are short and stable, making them one of the most precise clocks in the universe. As a result, pulsar binaries become unique laboratories to study stellar evolution (Ferdman et al., 2020; Stairs, 2004), NS magnetospheric physics (Breton et al., 2012; Lomiashvili & Lyutikov, 2014) and the strong gravitation near compact objects (Kramer et al., 2006; Ransom et al., 2014). Up to now, more than hundreds of pulsar binary systems have been found. However, among them, few binaries harbour another compact object besides pulsar. For instance, only fifteen double NS systems were detected, and thirteen of them were found to contain millisecond pulsars (Pol et al., 2019). Moreover, the pulsar-black hole (BH) binaries are even considered the holy grail of pulsar astronomy, but their only observed evidence so far is the gravitational-wave (GW) signals from those NS-BH merge events detected by the LIGO–Virgo detector network (Abbott et al., 2021; Collaboration et al., 2024).
@@
This picture is partly motivated by three recently found puzzling radio sources: GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839-10 (Hurley-Walker et al., 2023). Among them, PSR J0901-4046 (Caleb et al., 2022) shows variable pulse profiles and has the shortest period $\sim 1.3\,\rm min$, which is much longer than a normal pulsar. As for GLEAM-X J1627 (Hurley-Walker et al., 2022), its $\sim 18.2\,\text{min}$ period and multiple-peak pulse structure are similar to the observational appearance of a self-lensed pulsar-BH binary that our model predicts. GPM J1839-10 (Hurley-Walker et al., 2023) is also characterized by an ultra-long period of $\sim 22\,\text{min}$. Moreover, it has been found to have been actively existing for three decades, which makes this source even more mysterious.
This paper is organized as follows. In §2, we describe the toy model for the self-lensing and the assumed intensity profile of the pulsar’s intrinsic emission. We estimate the number of observable self-lensed pulsar-BH binaries in §3. In §4, we predict the appearance of ultra-long period radio signals in an edge-on pulsar-BH binary, apply the model to those recently found sources and check if they favour the self-lensing scenario as their physical origin. Additionally, we consider the Shapiro delay effect and estimate its resulting distortion of the lensed pulse profiles in §5. Finally, we summarize the results in §6.
@@ 2.2 Pulsar emission and self-lensing signals
With different parameter values, the lensed signals show diverse appearances. In Fig. 3, panels (a) and (c) exhibit the cases in which $M_{\rm BH}=10\rm M_{\odot}$. In comparison to (a), the sub-pulses in (c) have a higher level of $F_{0}$ and $\phi$, which makes the amplified $F_{0}$ comparable to the peak fluxes of the surrounding lensed sub-pulses and results in a larger duty cycle. For the cases in panels (b), (d) and (e), the BH mass is $20\rm M_{\odot}$. Among them, panel (b) showcases a situation where the pulsar spins faster than in (d) and (e), which enables more sub-pulses to be detected in a self-lensing signal. Therefore, the overall shape of the signal resembles the amplification factor evolution shown in Fig. 2. The orbital inclination is adjusted to higher ($cos\,i=3\times 10^{-4}$) and lower ($cos\,i=6\times 10^{-6}$) levels in panels (d) and (e), corresponding to flatter or steeper signal profiles, respectively. In addition, we present an extreme case in panel (f), where the BH mass $M_{\rm BH}=2\times 10^{4}\rm M_{\odot}$, the spin period $P=4\rm s$ and the orbital period $P_{\rm b}=0.33\rm hr$.
3 The number estimation for detection
Here we estimate the number of observable self-lensed pulsar-BH systems within the Milky Way. We first compute the number of galactic NS-BH binaries that can reach a threshold lensing amplification $A_{\rm th}$ in §3.1. This calculation also incorporates constraints from recent NS-BH merger events that LIGO-Virgo detects. Then in §3.2, we estimate the number of galactic self-lensed pulsar-BH binaries assuming a uniform radio luminosity for all pulsars.
3.1 The binary number estimation for a threshold amplification
$\tau=\frac{\pi R_{\rm E}^{2}}{4\pi a^{2}}\,=\,\frac{GM_{\rm BH}}{ac^{2}}.$ (10)
It is also the possibility for $u_{\rm min}\leq 1$, which corresponds to that the closest angular separation between the pulsar and BH is smaller than the angular Einstein radius. In observation, a threshold amplification $A_{\rm th}$ is needed to make the lensed pulse flux exceed the telescope’s sensitivity and become detectable (i.e., as long as $A_{\rm max}>A_{\rm th}$, signals composed of self-lensed pulses can be observed). This implies: $u_{\rm min}\leq u_{\rm th}=A_{\rm th}^{-1}$. The possibility that this constraint holds is:
$\tau(A_{\rm th})$ $=\,\frac{\pi(u_{\rm th}R_{\rm E})^{2}}{4\pi a^{2}}$ (11)
$p(M_{\rm BH})$ $\equiv\frac{1}{N_{\rm tot}}\frac{dN}{dM_{\rm BH}}$ (12)
where $t_{\rm coal}(P_{\rm b},M_{\rm BH})$ is the coalescence time given by Eq. (24), and $\mathcal{R}_{\rm i}$ is the merger rate, which varies depending on the binary type. Because the binary studied here is the NS-BH binary, $\mathcal{R}_{\rm i}=55\,\rm Gpc^{-3}\,yr^{-1}$ is adopted, based on the recent NS-BH coalescence events detected by LIGO-Virgo (Collaboration et al., 2024). Assuming that there are $10^{-2}$ Milky-Way-like galaxies per $\rm Mpc^{3}$ (Montero-Dorta & Prada, 2009), this number could be converted to $\mathcal{R}_{\rm i}\approx 5.5\,\rm MWEG^{-1}\,Myr^{-1}$, where the unit stand for the number of such mergers per Milky Way equivalent galaxy per Myr.
[…]
We set $M_{\rm NS}=1.4\,\rm M_{\odot}$ and allow $M_{\rm BH}$ to range from $M_{\rm BH,min}=10\,\rm M_{\odot}$ to $M_{\rm BH,max}=100\,\rm M_{\odot}$. To identify such systems, a sufficient number of signals must be observed. Therefore, we choose $P_{\rm min}=30\,\rm s$. This ensures that the binaries considered could all exist at least $\sim 1\rm yr$ until coalescence. Because the fraction of binaries with a short period ($\sim\rm min$) is small, $P_{\rm max}$ is set to be $\gg P_{\rm min}$. Thus Eq. (14) can be written as:
$\mathcal{N}_{\rm lens}(A_{\rm th})\approx 12\times\left(\frac{P_{\rm max}}{200\,\rm day}\right)^{2}\left(\frac{A_{\rm th}}{10}\right)^{-2}.$ (15)
This indicates that within the Milky Way, approximately 12 NS-BH binaries could satisfy the threshold amplification of $A_{\rm th}=10$.
3.2 The number of observable self-lensed pulsar-BH binaries
$f_{\rm c}=\frac{\Omega_{\rm c}}{4\pi}=\frac{1}{4}\,\sin^{2}(\phi).$ (16)
$\phi=\pi/12$ is set hereafter.
[…]
$\approx 0.7\left(\frac{P_{\rm max}}{200\,\rm day}\right)^{2}\left(\frac{A_{\rm th}}{2}\right)^{-2}\left(\frac{F_{\rm lim}}{10\,\rm mJy}\right)^{-1}\left(\frac{L_{\nu}}{1\,\rm Jy\,kpc^{2}}\right).$
Using $\rho(P_{\rm b},M)$ in a distance interval $dD$:
the number of the ultra-long period radio sources is:
[…]
$\approx 2.5\times\left(\frac{P_{\rm max}}{200\,\rm day}\right)^{2}\left(\frac{F_{\rm lim}}{10\,\rm mJy}\right)^{-1}\left(\frac{L_{\nu}}{1\,\rm Jy\,kpc^{2}}\right).$
In Figure 4, we explore the correlation between the estimated numbers ($\mathcal{N}_{\rm mod}$ and $\mathcal{N}_{\rm ulp}$) and different values of $P_{\rm max}$, $L_{\nu}$ and $F_{\rm lim}$. Notably, when $F_{\rm lim}$ is at a moderate level ($\gtrsim 0.1\,\rm mJy$ for black lines or $\gtrsim 1\,\rm mJy$ for red lines), these two numbers can be effectively described by the power-law function of the free parameters, as shown in the second lines of Eqs. (19) and (22). However, when the sensitivity reaches a level at which the telescope can detect the intrinsic pulsar emission from every pulsar-BH binary in the Milky Way, the observed signals will manifest as modulated pulsar signals, and no puzzling ultra-long period sources can be found. Consequently, $\mathcal{N}_{\rm mod}$ becomes steady and $\mathcal{N}_{\rm ulp}$ sharply drops.
As Eqs. (19) and (22) show, if all the pulsar’s radio luminosity is $1\,\rm Jy\,kpc^{2}$, there will be approximately $3$ pulsar-BH binaries in the Milky Way with observable self-lensing signals that can be detected by a telescope with a sensitivity of $10\,\rm mJy$. Among them, about $2$ binaries would exhibit ultra-long period characteristics and the rest are only modulated pulsar signals. The numbers $\mathcal{N}_{\rm mod}$ and $\mathcal{N}_{\rm ulp}$ are sensitive to $P_{\rm max}$, i.e., $\propto P_{\rm max}^{2}$, as shown in Eqs. (19) and (22). This suggests that the orbital periods of these binaries should be long, ranging from tens to hundreds of days.
@@ 4.1 Puzzling radio sources
When carrying out this work, we came across three recently discovered interesting radio sources, namely GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839-10 (Hurley-Walker et al., 2023). They all exhibit characteristics similar to extremely slow-rotating pulsars with spin periods of approximately $18.2\,\rm min$, $1.3\,\rm min$, and $22.0\,\rm min$, respectively.
@@
In contrast, GLEAM-X J1627 (Hurley-Walker et al., 2022) showcases signals with a much longer period ($18.2\,\rm min$) and a relatively stable pulse morphology, which exhibits a multi-peak structure. Its pulse duration ranges from $30$ to $60\,\rm s$, and the period of $18.18\,\rm min$ poses challenges to existing pulsar models. Even more puzzling, after three months of activity, it disappeared from observations and could not be found in available data. This ultra-long period but short-lived behaviour has prompted a range of interpretations, including a WD with relatively strong magnetic fields (Katz, 2022; Loeb & Maoz, 2022) and a spinning-down magnetar (Beniamini et al., 2023; Suvorov & Melatos, 2023; Tong, 2023).
@@
Therefore, we apply our model to these sources to check if they are actually self-lensed pulsars. In Fig. 5, PSR J0901-4046 is represented just by a blue dot due to its well-constrained pulse duration. For GLEAM-X J1627 and GPM J1839-10, we plot them as blue dots with error bars, where the dots correspond to the median value of their pulse durations and the bars represent a range between the maximum and minimum values. The results suggest BH masses of $4\,\rm M_{\odot}$ and $\sim 10^{4}\,\rm M_{\odot}$ for PSR J0901-4046 and GLEAM-X J1627, with corresponding coalescence times of $10^{2}\,\rm yrs$ and $\sim 6\times 10^{2}\,\rm yrs$, respectively. For GPM J1839-10, its variable signal duration makes the results range in several orders of magnitudes: The BH mass ranges from $\sim 10^{3}$ to $\sim 10^{6}\,\rm M_{\odot}$, with coalescence times spanning from $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$.
In our estimation, the coalescence time derived for GLEAM-X J1627 appears excessively long, inconsistent with its reported three-month active period. Conversely, PSR J0901-4046 has been staying active since its detection, which aligns well with the estimated 100-year lifetime. However, the derived mass for its lens is $\sim 4\,\rm M_{\odot}$, which lies in the $2-5\,\rm M_{\odot}$ mass gap and is not typical for observed stellar-mass BHs so far. Regarding GPM J1839-10, the derived coalescence time range ($\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$) aligns with its three-decade activity, implying monitoring a continuing activity of the system for at least another 30 years, which can be verified by the future monitoring.
@@
However, our model predicts negative, large period derivatives ($\sim 10^{-8}\,\rm s\,s^{-1}$), apparently inconsistent with the reported $\dot{P}$ of the three sources, especially for PSR J0901-4046: It has a positive $\dot{P}\approx 2.25\times 10^{-13}\,\rm s\,s^{-1}$ (Caleb et al., 2022). The other two sources, GLEAM-X J1627 and GPM J1839-10, have reported $\dot{P}$ below two upper limits: $\leq 1.2\times 10^{-9}\,\rm s\,s^{-1}$ (Hurley-Walker et al., 2022) and $\leq 3.6\times 10^{-13}\,\rm s\,s^{-1}$ (Hurley-Walker et al., 2023).
The above case applications of our self-lensing scenario to GLEAM-X J1627 and GPM J1839-10 suggest they are intermediate mass ratio inspirals (IMRIs), consisting of an NS and an intermediate-mass black hole (IMBH). One of the most promising environments to embed IMBHs is the centres of star clusters. Since the formation of an IMBH is related to the numerous low-mass compact objects that orbit around it, star clusters are also ideal places to find IMRIs.
Many works studied the formation and the evolving fate of IMRIs. Mandel et al. (2008) discussed four mechanisms that could form IMRIs, including the hardening of a CO-IMBH Binary through three-body interactions, Kozai resonance, direct capture via GW emission and inspiral of a CO from a tidally captured main-sequence star. Among them, the three-body interaction between an IMBH and a binary system is recognized as the primary pathway. Based on this, the estimated formation rate of the NS-IMBH IMRI systems that are detectable in GWs at the coalescence stage would be $\sim 0.1\,\rm Gal^{-1}\,Myr^{-1}$ (Mandel et al., 2008). This result was supported by N-body simulations (Konstantinidis et al., 2013; MacLeod et al., 2016; Fragione et al., 2018).
$M_{\rm BH}$ $t_{\rm coal}$ $\mathcal{R}_{\rm NSBH}$
$(\rm M_{\odot})$ $(\rm yrs)$ $\rm(MWEG^{-1}\,Myr^{-1})$
PSR J0901 $4$ 100100~{}100100 $10^{4}$
GLEAMX J1627 $10^{4}$ 600 $1.7\times 10^{3}$
GPM J1839 $10^{3-6}$ $3\times 10^{1-4}$ $3.3\times 10^{1-4}$
Table 1: The BH masses, coalescence times and merger rates for the three sources, derived by applying the self-lensing pulsar-BH scenario to them.
If these three sources were indeed self-lensed pulsar-BH binaries, we could roughly derive the underlying merger rates for each of them based on their estimated $t_{\rm coal}$ (see Table 1):
@@
$\mathcal{R}_{\rm NSBH}\approx 10^{4}\times\left(\frac{t_{\rm coal}}{100\,\rm yrs}\right)^{-1}\,\rm MWEG^{-1}\,Myr^{-1},$ (25)
where $\rm MWEG$ stands for Milky Way equivalent galaxies, as these sources are found within the Milky Way. On the other hand, the LIGO-Virgo detection of NS-BH mergers gives a merger rate estimate of $\mathcal{R}_{\rm i}=5.5\,\rm MWEG^{-1}\,yr^{-1}$ (see §3.1), which is much lower than $\mathcal{R}_{\rm NSBH}$, especially for PSR J0901-4046.
Note that for GLEAM-X J1627 and GPM J1839-10, their corresponding $\mathcal{R}_{\rm NSBH}$ cannot be directly compared to $\mathcal{R}_{\rm i}$, since the GWs from a NS’s inspiral into an IMBH is currently difficult to detect. However, as mentioned above, the anticipated event rate of GW-observable NS-IMBH IMRI is just $\sim 0.1\,\rm Gal^{-1}\,Myr^{-1}$ (Mandel et al., 2008). Additionally, Fragione et al. (2018) estimated that there would be $\sim 10^{3}$ IMBHs within $1\,\rm kpc$ from the Galactic centre, and the IMBH-SBH merger rate in the local universe is  $\sim\,0.1-1\,\rm MWEG^{-1}\,Myr^{-1}$. Though NSs may be more common than SBHs, it is hard to imagine that the corresponding merger rate could be higher by several orders of magnitude.
Therefore, we conclude that these three puzzling radio sources are unlikely the self-lensed pulsars, because the predicted period decay rates exceed the observational constraints, and the implied merger rates are too high.